OpenStudy (anonymous):

A solution is prepared by mixing 90.0 mL 0.439 M NaCl, 100.0 mL 0.0516 M MgCl2, and 235.0 mL of water. What are the molarities of [Na+], [Mg2+], and [Cl -] in the resulting solution? Please tell me how I would go about solving this.

OpenStudy (anonymous):

1) write an equation for the possible reaction 2) determine based on solubility rules if a reaction will occur. meaning if a precipitate will form

OpenStudy (anonymous):

It doesn't seem like to me that a reaction will occur, so how do I find the molarities of each individual element?

OpenStudy (anonymous):

Do you know the formula for molarity? Can you calculate the number of moles of Na in the original solution of 90ml of 0.439M

OpenStudy (anonymous):

Yea, moles solute over liters solution, but Na is bonded with Cl, how do I find Na and not NaCl?

OpenStudy (anonymous):

Based on solubility rules what will happen to NaCl in water?

OpenStudy (anonymous):

It becomes Na + Cl, right? So do I just used the values given to me for NaCl to find Na?

OpenStudy (anonymous):

Or do I convert NaCl to Na?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

For every mole of NaCl there will be one mole of Na

OpenStudy (anonymous):

Oh! Since it's moles, we can just say the values are the same?

OpenStudy (anonymous):

Correct

OpenStudy (anonymous):

So would I just say that the molarity of Na is the same as the given molarity of NaCl?

OpenStudy (anonymous):

The original molarity of Na is the same as the original molarity of NaCl. However, the molarity changes; since M = moles solute / L solution, the volume of the original molarity changes since MgCl2 and water is being added.

OpenStudy (anonymous):

First calculate the number of moles Na in the original NaCl solution. Please post your answer.

OpenStudy (anonymous):

439 moles per 1 liter? And do I find what it is in 90 mL then?

OpenStudy (anonymous):

You meant 0.439 mole per L. Yes find what it is in 90 ml. What do you need to convert ml to?

OpenStudy (anonymous):

Liters, so it would be something like (0.439 mol Na) x (90mL/1L)?

OpenStudy (anonymous):

I'm not sure what you mean. There are two methods of doing it: Method 1) \[\frac{ ml }{ 1 }*\frac{ L }{ ml }*\frac{ mol }{ L }\]

OpenStudy (anonymous):

Method 2: convert ml to L \[M = \frac{ mole solute }{ L solution }\] Plug in the value for M and L and algebraically solve for moles

OpenStudy (anonymous):

Ok, my teacher likes dimensional analysis/factor labelling/whatever you want to call it, so I would use method 1. I got 0.03951 moles as my answer.

OpenStudy (anonymous):

Although that should be 0.040 moles with sig figs.

OpenStudy (anonymous):

Correct. So now calculate the new molarity, using 0.03951 moles / (original solution + MgCl2 solution + water)

OpenStudy (anonymous):

Keep the 0.03951 in your calculator round to sig figs in your last step/

OpenStudy (anonymous):

I would love to, but my teacher tells us to do sig figs for every step, so I'll do it her way.

OpenStudy (anonymous):

That's fine with me

OpenStudy (anonymous):

So 0.040 moles / 425mL of everything added together is 0.040 moles / 0.425 L of total volume? So 0.094 M?

OpenStudy (anonymous):

Your calculation is correct; you are missing a sig fig with 0.094 M.

OpenStudy (anonymous):

Oh yea, true. So it's 0.0941.

OpenStudy (anonymous):

Correct. You can do the same now for Mg2+ and Cl-. Just be sure to use the mole Cl from NaCl (which is a 1:1 ratio) and mol Cl- from MgCl2 which is a 1:2 mole ratio of 1 mol MgCl2 : 2 mol Cl-. Good Luck

OpenStudy (anonymous):

is this high school chemistry or college chemistry?

OpenStudy (anonymous):

High school, but accelerated.

OpenStudy (anonymous):

One thing: For Cl do I add the moles and liters of it from NaCl and MgCl2 and do a total?

OpenStudy (anonymous):

You'll have a much easier in time in college compared to your classmates that did not have accelerated chemistry.

OpenStudy (anonymous):

I'm glad to hear it, this is confusing so far.

OpenStudy (anonymous):

\[\frac{ (molCl^- fromNaCl)+(molCl^- from MgCl2) }{ L soltn NaCl2 + L soltn MgCl2 + L water }\]

OpenStudy (anonymous):

Ok, that makes sense. Thanks for helping, I'll get around to this soon.

OpenStudy (anonymous):

The more you practice the less confusing it gets. It starts to make sense a semester to a year later.

OpenStudy (anonymous):

I hope so. Ok, going to go eat and then try this after.